OpenStudy (anonymous):

Hi, I really need help. I was on a homework help site for 140 minutes, and as soon as it was my turn, it logged me out. I need help with Ohm's Law, I understand nothing. I am in grade nine. Please help me. First of all, what is Ohm's Law? (Science, Physics)

jimthompson5910 (jim_thompson5910):

I'm not that great at physics or electrical engineering but according to this page http://hyperphysics.phy-astr.gsu.edu/hbase/electric/ohmlaw.html Ohms law is I = V/R I = current V = voltage R = resistance

OpenStudy (anonymous):

@jim_thompson5910 thanks! I kind of understand that, but my teacher taught us like 10 equations, and I know nothing and I have a unit test tomorrow. Thanks for looking that up though

jimthompson5910 (jim_thompson5910):

which 10 equations are you supposed to know? basically equations like that?

OpenStudy (anonymous):

I really don't know. we did some equations involving circuits, and then we did some about phantom load. like when the tv is plugged in for 3 weeks, how much energy is used

OpenStudy (anonymous):

Jim already explained it...Ohm’s law is electric current=voltage/resistance

OpenStudy (anonymous):

What part you don't understand

OpenStudy (anonymous):

I understand that, but my teacher taught us a bunch of different equations related to it

OpenStudy (anonymous):

i will find one of my questions and post it just a sec

thomaster (thomaster):

\(\large\Omega=\frac{V}{A}\) is also one :D

OpenStudy (anonymous):

Look at this website...it may help you http://hyperphysics.phy-astr.gsu.edu/hbase/electric/ohmlaw.html

OpenStudy (anonymous):

Anyway, I'll wait for the question to be posted

OpenStudy (anonymous):

Thanks for adding that thomaster

OpenStudy (anonymous):

@blast234 thanks! I am kind of dumb when it comes to physics...so whats a conducter

OpenStudy (anonymous):

just finding it one sec

jimthompson5910 (jim_thompson5910):

something that allows electricity or electric charge to flow through it contrast that with an insulator which does not let charge flow through

OpenStudy (anonymous):

a conductor conducts electricty

hero (hero):

I know all of the equations

OpenStudy (anonymous):

Conductors are materials that permit electrons to flow freely from atom to atom and molecule to molecule.

OpenStudy (anonymous):

oh my goodness. i cant stand physics. there are so many things i dont know. like... percent efficiency... Eout over Ein x100%

OpenStudy (anonymous):

ok... i tthink i get that. like in a light bulb. how much energy goes in, how much comes out and the percent is the waste... or the perecent that actually comes out?

jimthompson5910 (jim_thompson5910):

that's like saying if you put in 20 units of effort into the system and get 10 units of effort back out, then you're getting (Eout)/(Ein)*100 = (10/20)*100 = (1/2)*100 = 0.5*100 = 50% efficiency

jimthompson5910 (jim_thompson5910):

so yeah, it helps you see how much is being wasted, in this case 50% is wasted

OpenStudy (anonymous):

Efficiency is measured by Work out/Energy in . Heat is not work. Work + heat always amounts to 100% of the energy.

OpenStudy (anonymous):

what is E=Pxt

OpenStudy (anonymous):

and C= ExR???

OpenStudy (anonymous):

IM GUESSING THIS HAS NOTHING TO DO WITH OHM'S LAW ,,,

OpenStudy (anonymous):

Energy = Power x Time. is E=Pxt

OpenStudy (anonymous):

2. A microwave was used 5 hours a day for 7 days. This power required for this microwave is 2.5 kW. Cost much would it cost to run the microwave for 1 week. The rate for the usage of electricity is $0.07 kWh (4pts)

OpenStudy (anonymous):

this is Eout over Ein x100% right?

jimthompson5910 (jim_thompson5910):

no you just do a bunch of conversions really 5 hours/day * 7 days = 35 hours total 35 hrs * 2.5 kW = 87.5 kWhours (87.5 kWhours)*(0.07 dollars/1kWhour) = 6.125 = 6.13 dollars

OpenStudy (anonymous):

so would this be,,, 5x2.5kW= 12.5kW 12.5 kWx7 = 87.5kWh 87.5x the cost???

OpenStudy (anonymous):

Oh thanks!!

jimthompson5910 (jim_thompson5910):

yeah it uses up 87.5 kW total (over 35 hours), you then use that to find the total cost

OpenStudy (anonymous):

how do i convert the watts? because i remember my teacher saying something about kilowatts, mega watts and other stuff

jimthompson5910 (jim_thompson5910):

1000 watts = 1 kilowatt

OpenStudy (anonymous):

are u guys in the same grade or in a higher grade than me?

OpenStudy (anonymous):

oh right right! i need to remember that!

jimthompson5910 (jim_thompson5910):

1,000,000 watts = 1 megawatt

OpenStudy (anonymous):

hopefully she'll put that in the test for our class to remember

OpenStudy (anonymous):

Explain the difference between how a non renewable recourse can create electricity vs a renewable resource.

OpenStudy (anonymous):

what is this?

jimthompson5910 (jim_thompson5910):

just remember the prefixes and you'll be fine kilo - thousand watt - watt so kilo + watt = thousand watts

OpenStudy (anonymous):

what is a watt? the unit to measure electricity?

jimthompson5910 (jim_thompson5910):

it's a measure of energy consumption I think

OpenStudy (anonymous):

i really hope ur in a higher grade than me... or else i'll be very sad lol

OpenStudy (anonymous):

ok...thanks

jimthompson5910 (jim_thompson5910):

oh to be more precise, it's a unit of power where power = energy/time

OpenStudy (anonymous):

thanks so much this really means alot

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

r u in a higher grade than me?

jimthompson5910 (jim_thompson5910):

tbh, you're better off asking this in either the science, physics, computer science sections since this really isn't math you'll find better experts in there

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

wait what? omg. i havnt been on here since january, and i never logged out. i guess i was in a big rush, and i never checked!!!

OpenStudy (anonymous):

do u have instragram. u dont have to reply if u dont want to lol im not tryna be a stalker

jimthompson5910 (jim_thompson5910):

no have never used the service

jimthompson5910 (jim_thompson5910):

and yes you should check out the other sections to see if they have what you're looking for

OpenStudy (anonymous):

oh ok. thanks a lot!!!1

OpenStudy (anonymous):

thanks u so much! bye now!

jimthompson5910 (jim_thompson5910):

yw, good luck

OpenStudy (anonymous):

@jim_thompson5910 thanks im going to need it. the unit test is tomorrow and i crammed. hopefully, though, i'll do beeter on the final exam mid june. im going to start styding now

thomaster (thomaster):

do you use a graphing calculator? If you do you can use a program where you can insert all the formulas. I did this for my physics test a few years ago the teachers even recomended it. It takes away much stress if you don't have to remember a thousand formulas which you aren't going to remember after the test anyway.